how do you solve 16 m^2 = ____cm^2
Found 2 solutions by lynnlo, ikleyn:
Answer by lynnlo(4176)
Answer by ikleyn(53742)

.

1 meter = 100 centimeters.


1 square meter = 100^2 cm^2 = 10,000 cm^2.


16 m^2 = 16*10,000 = 160,000 cm^2.



ANSWER.  16 m^2 = 160,000 cm^2.

It is not simple - - - it is SUPER-simple conversion.